OpenStudy (anonymous):

3. Find the circumcenter of triangle ABC with A(1, 6) anf C(5, 4). (1 point)

OpenStudy (anonymous):

I think you forget to include the third point

OpenStudy (anonymous):

oops b(1, 4)

OpenStudy (anonymous):

find the line bisector for two sides

OpenStudy (anonymous):

in addition find a slope for the two lines

OpenStudy (anonymous):

The line passing through each midpoint is perpendicular to the line joining two vertices.

OpenStudy (anonymous):

I am sure you know the relation between the slopes of two perpendicularly intersected lines.

OpenStudy (anonymous):

hmm. i kinda lost cuz i was really sick for a few weeks, so to get back on track i just skipped through the stuff and didnt really learn it.

OpenStudy (anonymous):

well, the midpoint between two points (x1,y1) and (x2,y2) is \[(\frac{ x1+x2 }{ 2 },\frac{ y1+y2 }{ 2 })\]

OpenStudy (anonymous):

would you calculate the midpoint between A and B so that I make sure you understand it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

1 + 2/2 6 + 8/2 ?? am i right so far

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Those numbers you see next to x and y are subscripts not multipliers

OpenStudy (anonymous):

\[(\frac{ 1+1 }{ 2 },\frac{ 4+6 }{ 2 })\])

OpenStudy (anonymous):

yeah i didnt think so but i wasnt sure, 1,1